# Dynamic On-Palm Manipulation via Controlled Sliding

### v1

@@ I Introduction
Recent advancements in robot manipulation have demonstrated impressive dexterity [6] and generality [10] [9]. However, these methods largely focus on slow tasks that can be viewed from a quasi-static perspective. As robots are increasingly being asked to perform manipulation tasks in logistics applications such as warehouse robotics, speed becomes a key driving metric. While there are plenty of examples of dynamic manipulation, the methods are often achieved using ad-hoc, task-specific solutions [33] and demonstrated on systems with few degrees of freedom and few contacts. The desire for a general control framework for contact-rich tasks has resulted in many formulations for contact-implicit model predictive controllers [2] [24] [25] (MPC), which can automatically plan contact, when and where to make and break contact, and are reported to be fast enough for real-time control.
@@
  • Demonstrate the generality of the contact-implicit MPC framework with a task where the robot must rotate the tray using a wall.
@@ IV-A Linear Complementarity Model
We model the dynamics of the end effector, object, and external contacts as a discrete time LCS. We ignore rest of the robot arm in the MPC model by considering the end effector as an isolated floating object with only translation degrees of freedom and controller directly with forces applied to its center of mass. To ensure downstream feasibility when applying this model to the actual system, we impose workspace and input limits on the MPC model. The state of the LCS $x_{lcs}=[q_{lcs},v_{lcs}]$ is a combination of the positions $q_{lcs}=[q_{ee},q_{obj}]$ of the end effector and object and the corresponding velocities $v_{lcs}=[v_{ee},v_{obj}]$. The control input $u_{lcs}$ to the LCS are forces applied directly to the end effector center of mass, such that it can be controlled in 3D translation. Finally, the contact forces $\lambda$ are the interaction forces between the end effector, object, and external supports.
@@
where $x_{k}\in\mathbb{R}^{n_{x}}$, $\lambda_{k}\in\mathbb{R}^{n_{\lambda}}$, and $u_{k}\in\mathbb{R}^{n_{u}}$. Eq. 1 are the system dynamics linearized at the current state and input, $x_{lcs},u_{lcs}$. The $\perp$ indicates a complementarity constraint, where $0\leq\lambda\perp\phi\geq 0$ implies $\lambda\geq 0$, $\phi\geq 0$, $\lambda^{T}\phi=0$. Critically, this constraint succinctly describes the multi-modality of contact dynamics for both making and breaking contact as well as the boundary between stick and slip. For example, the boundary between sliding and sticking friction for a given sliding direction can be expressed as:
@@
where $\lambda_{n}$ is the normal force, and $\lambda_{t}$ is the tangential force in the opposite direction of the sliding velocity $v$. With this context, Eq. 2 is the linearization of the contact boundaries with respect to the current state and input.
The contact dynamics of the LCS (Eq. 2) are governed by our choice of contact geometry. We approximate the surface-surface contacts between the end effector and object as well as the object and external contacts using point contacts. We use three contact points between the end effector and object, because that is the minimum number necessary to have statically stable surface-surface contact. Similarly, we model each support as two points to represent the line contacts. These contact geometries are visualized in Fig. 4. Under this choice of contact geometry, $\phi$ encodes the distance between any of these contact points and the tray, represented as a cylinder. Additionally, the number of contacts $n_{\lambda}$ is fixed. Note, we are ignoring potential contacts between the end effector and the supports. As these contacts are undesirable, we simply avoid these contacts by imposing workspace constraints on the end effector.
Figure 4: We consider seven total contacts for our task. The contact geometries shown in red. We represent the tray as a flat cylinder and we choose fixed contact points on the end effector and supports, which we model as spheres. The radii for the spheres are enlarged by a factor of 10 for visibility purposes. A minimum of three contact points are required to approximate surface-surface contact between the end effector and tray, while two contact points are required to model each line contact from the supports.
@@ V-A Complementary Consensus Control
where $N$ is the planning horizon, $Q_{f},Q,R$ are cost matrices, and $\mathcal{C}$ is a convex set representing state and input constraints. Eq. 6, Eq. 7, and Eq. 8 are the dynamics constraints of the LCS. Eq. 9 is the orthogonality constraint for the complementarity. Note, Eq. 9 is non-convex, but it is possible to introduce binary variables to represent the contact modes and transcribe the entire problem as a Mixed Integer Quadratic Program (MIQP). However, this scales poorly with the number of contacts, as a binary variable is needed for each contact across all knot points.
Instead, we adopt a method called Complementarity Consensus Control (C3) [2], which addresses the scaling problem by decoupling the time dependence of the contact decisions. The algorithm is based in consensus alternating direction method of multipliers (ADMM), which optimizes over multiple copies of the decision variables. The full details of the algorithm is outside of the scope of this paper, but a key property is that the algorithm alternates between solving the MPC problem as a quadratic program (QP) without complementarity constraints and projecting the current MPC solution to the complementarity constraints as a mixed integer quadratic program (MIQP) separately for each knot point. While the solutions will eventually converge to each other, we choose to terminate early after a fixed number of iterations on a suboptimal solution. We choose to terminate after the QP step, because we empirically observe better performance. Note, the suboptimal solutions from terminating early do not necessarily satisfy the full LCS dynamics, but in practice are good enough when used in MPC for even contact-rich tasks.
@@ V-B System Linearization
We approximate our system as a LCS at each C3 iteration. The continuous dynamics parameters ($A$, $B$, $d$) of the LCS can be solved via automatic differentiation using any popular rigid body dynamics library. The gap function $\phi$ and corresponding contact Jacobians $J$ for convex geometries can be computed by a library that implements collision detection, e.g. via the GJK algorithm[13]. With $\phi$ and $J$ for each contact along with a choice of force basis, we can compute the contact-related terms: $D$, $E$, $F$, $H$, $c$. We use the convex time-stepping contact model proposed by Anitescu and Potra [1] to form the force basis. In this model, contact forces are parameterized via the extreme rays of the pyramidal approximation of the friction cone. That is, for a square pyramidal approximation, there are 4 contact force variables per point contact. This choice of contact force basis is visualized in Fig. 7.
@@ V-C MPC Modifications for Dynamic Motions
Warm starting by giving the MPC an initial guess from the previous solve is a common technique to reduce computation time. However, the values from the previous solution are often poor initial guesses because the contact modes at each knot point planned from the current MPC state may differ from the previous solution. Because the dynamics can vary greatly between contact modes, so can the values for $x,u,\lambda$. For this reason, we use the corresponding predicted values from the previous solution when possible for warm starting. Additionally, C3 involves multiple QP and MIQP solves per MPC solve each with different cost parameters. We address this by treating each QP and MIQP as separate optimization programs each with a separately cached set of warm start variables in order to increase the chance of good initial guesses.
@@ V-D Operational Space Control
s.t. $M\ddot{q}+C=g+u+J_{\lambda}^{T}\lambda$ (15)
@@
where $i$ refers to each task space objective and whose acceleration $\ddot{y}_{i}$ is linear mapping of $\ddot{q}$ and can be derived from the task space kinematics function $\phi_{i}$, where $y=\phi_{i}(q)$. Differentiating with respect to time, we have $\dot{y}=\frac{\partial\phi}{\partial q}\dot{q}=J_{y,i}\dot{q}$, and differentiating once more we get $\ddot{y}=\dot{J}_{y,i}\dot{q}+J_{y,i}\ddot{q}$. Eq. 15 is the dynamics constraint that relates actuator inputs $u$ to joint accelerations $\ddot{q}$ as explained in Eq. 4 and the actuator lower and upper limits are defined by $u_{min}$ and $u_{max}$ respectively.
@@ V-D1 End Effector Force Target
The end effector force target is an important component to accurately tracking the MPC plan without relying on overly stiff impedance gains or an integral term, both of which could cause instability for this task. To see this, consider the scenario where the robot balances the tray. Without a force target, the robot will not compensate for the weight of the object, and the object will sag according to the impedance stiffness $K_{p}$. While tracking error for interactions solely between the manipulator and object scales with stiffness, tracking error for systems with additional contacts is more complex. For example during the sliding maneuver, even small forces applied by the end effector to the object can result in significant effects on the weight distribution of the object across the supports and end effector. Because our task is governed by controlling friction forces, the weight distribution is important to accurately track.
@@ VI Experiments
Figure 5: System diagram for the hardware implementation. The different colored boxes indicate separate processes which a connected via arrows that indicate represent communication via ROS/LCM.
@@ VI-A Task Parameters
The exact position targets are given in Table I. The positions are defined in the world frame where the robot base is at the origin. For the first and second targets, the end effector target is the same as the tray, just offset in the vertical position to compensate for the thicknesses. Critically, the third target for the tray is outside the workspace limits of the end effector, so for this pair of targets the end effector target is chosen as the closest point within the feasible workspace. We choose these targets in order to be able to repeatedly execute the experiment without manually resetting the task environment. As detailed in Section III, the next target is only given when the tray reaches the previous target. We define reaching the target as being within 5 cm from the target location.
@@ VI-A1 Tray and End Effector
We use a standard circular food service tray with a smooth low friction bottom surface and a rubberized high friction upper surface and we model the tray simply as a cylinder with uniform density. For the end effector, we use a thin circular aluminum disk. Because the coefficient of friction between the machined aluminum and the tray’s bottom surface is not sufficiently high, we cover the top surface of the end effector with tape. We estimate the friction coefficients by slowly tilting the supports or end effector until the tray slips and using that angle to determine a single friction coefficient, assuming that the static and dynamic coefficients are identical. Detailed parameters for both objects are listed in Table II and the objects are shown in Fig. 6.
@@ VI-B Implementation
Both C3 and the OSC were implemented in C++ with the help of the robotics toolbox [38]. All code will be released, open-source, following the double-blind peer review period. Both controllers, as well as franka ros and the LCM to ROS bridges, are run on the same desktop with a Intel i7-8700K processor. The QP step of C3 was solved using OSQP [35], while the MIQP projection was solved using Gurobi [14]. The OSC QP was solved using OSQP [35] at 1000 Hz. We tune the OSC and C3 parameters by executing the task in the Drake [38] simulator using the hydroelastic contact model [28]. We directly apply the parameters that were tuned in simulation on the hardware implementation.
@@ VII Results
We performed multiple experiments to validate the robustness and generality of our framework. First we ablate our design decision to include the force tracking objective in the OSC by running experiments with and without that objective. We then demonstrate the reliability by continuously executing the experiment without manual resetting. Then, we demonstrate the robustness of our method to inaccurate models of mass and inertia by placing objects on the tray. We use the same controller parameters for all targets and across all the demonstrations, and footage of the experiments can be seen in the supplemental video.
@@ VII-A Force Tracking Ablation
First we ablate the contribution of tracking the end effector force by executing 10 experiments with and without the tracking objective. The tracking controller with the end effector force objective succeeded for 80% of the trials, failing once when trying to reach the second target when lifting an unbalanced tray and failing once to reach the third target when the tray slipped off in the direction of the robot base. The trajectory of the end effector and tray for one of the successful executions is shown in Fig. 8. The tracking controller without the end effector force objective succeeded for 30% of the trials, failing seven times to reach the third target due to the tray either colliding when the supports or slipping off in the direction of the robot base.
@@ VII-B Reliability Test
The reliability of our method is also evaluated by repeatedly executing the task without intervention - initializing the robot from its state after it successfully reaches the third target. This is possible without manual resetting because the final target position coincides with the initial position, and thus we treat tracking error from the previous execution as unstructured perturbations to the initial state. Our method was able to complete six full cycles before failing due to the tray reach the position threshold (within 5cm) of the third target.
@@ VIII Limitations
While our controller is fairly robust to mass and inertia, it is not robust to the height of the supports. This is not surprising as the tray, supports, and robot are stiff, leading to senstivity at the boundary between contact and no-contact. However, this can be addressed by quickly updating the LCS parameters [22] or generating the contact geometry and system dynamics entirely [20] [30].
Another limitation of this method originates from the LCS model used by the MPC. A known limitation of the LCS model is that it only has a linear representation of the contact decisions, meaning that it cannot consider contacts beyond the contact boundary. For example, if the end effector moves from underneath the tray, the gap function $\phi$ will shift and instead consider only horizontal contact between the end effector and tray. In this scenario, the MPC cannot find a solution for the task as it cannot reason about a path to go back under the tray.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1845298.

### v2

@@ I Introduction
Recent advancements in robot manipulation have demonstrated impressive dexterity [6] and generality [10] [9]. However, these methods largely focus on slow tasks that can be viewed from a quasi-static perspective. As robots are increasingly being asked to perform manipulation tasks in logistics applications such as warehouse robotics, speed becomes a key driving metric. While there are plenty of examples of dynamic manipulation, the methods are often achieved using ad-hoc, task-specific solutions [33] and demonstrated on systems with few degrees of freedom and few contacts. The desire for a general control framework for contact-rich tasks has resulted in many formulations for contact-implicit model predictive control [2] [24] [25] (MPC), which can automatically plan when and where to make and break contact and are reported to be fast enough for real-time control.
@@
  • Demonstration of the generality of the contact-implicit MPC framework with a task where the robot must rotate a circular tray using a wall.
@@ IV-A Linear Complementarity Model
We model the dynamics of the end effector, object, and external contacts as a discrete time LCS. We ignore rest of the robot arm in the MPC model by considering the end effector as an isolated floating object with only translation degrees of freedom and controlled directly with forces applied to its center of mass. To ensure downstream feasibility when applying this model to the actual system, we impose workspace and input limits on the MPC model. The state of the LCS $x_{lcs}=[q_{lcs},v_{lcs}]$ is a combination of the positions $q_{lcs}=[q_{ee},q_{obj}]$ of the end effector and object and the corresponding velocities $v_{lcs}=[v_{ee},v_{obj}]$. The control input $u_{lcs}$ to the LCS are forces applied directly to the end effector center of mass, such that it can be controlled in 3D translation. Finally, the contact forces $\lambda$ are the interaction forces between the end effector, object, and external supports.
@@
where $x_{k}\in\mathbb{R}^{n_{x}}$, $\lambda_{k}\in\mathbb{R}^{n_{\lambda}}$, and $u_{k}\in\mathbb{R}^{n_{u}}$ are the state, force, and input variables at the $k$-th knot point. Eq. 1 are the system dynamics linearized at the current state and input. The $\perp$ indicates a complementarity constraint, where $0\leq\lambda\perp\phi\geq 0$ implies $\lambda\geq 0$, $\phi\geq 0$, $\lambda^{T}\phi=0$. Critically, this constraint succinctly describes the multi-modality of contact dynamics for both making and breaking contact as well as the boundary between stick and slip. For example, the boundary between sliding and sticking friction for a given sliding direction can be expressed as:
@@
where $\lambda_{n}$ is the normal force, and $\lambda_{t}$ is the tangential force in the opposite direction of the sliding velocity $v$. With this context, Eq. 2 is the linearization of the contact boundaries at the current state and input.
The contact dynamics of the LCS are governed by our choice of contact geometry. We approximate the surface-surface contacts between the end effector and object as well as the object and external contacts using point contacts. We use three contact points between the end effector and object, because that is the minimum number necessary to have statically stable surface-surface contact. Similarly, we model each support as two points to represent the line contacts. These contact geometries are visualized in Fig. 4. Under this choice of contact geometry, $\phi$ encodes the distance between any of these contact points and the tray, represented as a cylinder. Under this modeling choice, the number of contacts $n_{\lambda}$ is fixed. Note, we are ignoring potential contacts between the end effector and the supports. As these contacts are undesirable, we simply avoid these contacts by imposing workspace constraints on the end effector.
Figure 4: We consider seven total contacts for our task. The contact geometries shown in red. We represent the tray as a cylinder and we choose fixed contact points on the end effector and supports, which we model as spheres. The radii for the spheres are enlarged by a factor of 10 for visibility purposes. A minimum of three contact points are required to approximate surface-surface contact between the end effector and tray, while two contact points are required to model each line contact from the supports.
@@ V-A Complementary Consensus Control
where $N$ is the planning horizon, $Q_{f},Q,R$ are cost matrices, and $x_{min},x_{max},u_{min},u_{max}$ are bounds on the state and input variables. Eq. 6, Eq. 7, and Eq. 8 are the dynamics constraints of the LCS. Eq. 9 is the orthogonality constraint for the complementarity. Note, Eq. 9 is non-convex, but it is possible to introduce binary variables to represent the contact modes and transcribe the entire problem as a Mixed Integer Quadratic Program (MIQP). However, this scales poorly with the number of contacts, as a binary variable is needed for each contact across all knot points.
Instead, we adopt a method called Complementarity Consensus Control (C3) [2], which addresses the scaling problem by decoupling the time dependence of the contact decisions. The algorithm is based in consensus alternating direction method of multipliers (ADMM), which optimizes over multiple copies of the decision variables. The full details of the algorithm is outside of the scope of this paper, but a key property is that the algorithm alternates between solving the MPC problem as a quadratic program (QP) without complementarity constraints and projecting the current MPC solution to the complementarity constraints as a mixed integer quadratic program (MIQP) separately for each knot point. While the solutions will eventually converge to each other, we choose to terminate early after a fixed number of iterations on a potentially suboptimal solution. We choose to terminate after the QP step, because we empirically observe better performance. Note, the suboptimal solutions from terminating early do not necessarily satisfy the full LCS dynamics, but in practice are good enough when used in MPC for even contact-rich tasks.
@@ V-B System Linearization
We approximate our system as a LCS at each C3 solve. The continuous dynamics parameters ($A$, $B$, $d$) of the LCS can be solved via automatic differentiation using any popular rigid body dynamics library. The gap function $\phi$ and corresponding contact Jacobians $J$ for convex geometries can be computed by a library that implements collision detection, e.g. via the GJK algorithm[13]. With $\phi$ and $J$ for each contact along with a choice of force basis, we can compute the contact-related terms: $D$, $E$, $F$, $H$, $c$. We use the convex time-stepping contact model proposed by Anitescu and Potra [1] to form the force basis. In this model, contact forces are parameterized via the extreme rays of the pyramidal approximation of the friction cone. That is, for a square pyramidal approximation, there are 4 contact force variables per point contact. This choice of contact force basis is visualized in Fig. 7.
@@ V-C MPC Modifications for Dynamic Motions
Warm starting by giving the MPC an initial guess from the previous solve is a common technique to reduce computation time. However, the values from the previous solution are often poor initial guesses because the contact modes at each knot point planned from the current MPC state may differ from the previous solution. Because the dynamics can vary greatly between contact modes, so can the values for $x,u,\lambda$. For this reason, we use the corresponding predicted values from the previous solution when possible for warm starting. Additionally, C3 involves multiple QP and MIQP solves per MPC solve each with different cost parameters. We address this by treating each QP and MIQP as separate optimization programs each with a separately cached set of warm start variables in order to increase the quality of the warm start.
@@ V-D Operational Space Control
s.t. $M\ddot{q}+C=u+J_{\lambda}^{T}\lambda$ (15)
@@
where $i$ refers to each task space objective and whose acceleration $\ddot{y}_{i}$ is linear mapping of $\ddot{q}$ and can be derived from the task space kinematics function $\psi_{i}$, where $y=\psi_{i}(q)$. Differentiating with respect to time, we have $\dot{y}=\frac{\partial\psi}{\partial q}\dot{q}=J_{y,i}\dot{q}$, and differentiating once more we get $\ddot{y}=\dot{J}_{y,i}\dot{q}+J_{y,i}\ddot{q}$. Eq. 15 is the dynamics constraint that relates actuator inputs $u$ to joint accelerations $\ddot{q}$ and the actuator lower and upper limits are specified by $u_{min}$ and $u_{max}$ respectively.
@@ V-D1 End Effector Force Target
The end effector force target is an important component to accurately tracking the MPC plan without relying on overly stiff impedance gains or an integral term, both of which could cause instability for this task. To see this, consider the scenario where the robot balances the tray. Without a force target, the robot will not compensate for the weight of the object, and the object will sag according to the impedance stiffness $K_{p}$. While tracking error for interactions solely between the manipulator and object scales with stiffness, tracking error for systems with additional contacts is more complex. For example during the sliding maneuver, even small forces applied by the end effector to the object can result in significant effects on the weight distribution of the object across the supports and end effector. Because our task is governed by friction forces, this objective is particularly important.
@@ VI Experiments
Figure 5: System diagram for the hardware implementation. The different colored boxes indicate separate processes which are connected via arrows that indicate represent communication via ROS/LCM.
@@ VI-A Task Parameters
The exact position targets are given in Table I. The positions are defined in the world frame where the robot base is at the origin. For the first and second targets, the end effector target is the same as the tray, just offset in the vertical position to compensate for the thicknesses. Critically, the third target for the tray is outside the workspace limits of the end effector, so for this pair of targets the end effector target is chosen as the closest point within the feasible workspace. We choose to make the third target to be identical to the initial position in order to be able to repeatedly execute the experiment without manually resetting the task environment. As detailed in Section III, the next target is only given when the tray reaches the previous target. We define reaching the target as being within 5 cm from the target location.
@@ VI-A1 Tray and End Effector
We use a standard circular food service tray with a smooth low friction bottom surface and a rubberized high friction upper surface. We model the tray as a cylinder with uniform density. We machine the disk-shaped end effector out of aluminum. Because the coefficient of friction between the machined aluminum and the tray’s bottom surface is not sufficiently high, we cover the top surface of the end effector with tape. We estimate the friction coefficients by slowly tilting the supports or end effector until the tray slips and using that angle to determine a single friction coefficient, assuming that the static and dynamic coefficients are identical. Detailed parameters for both objects are listed in Table II and the objects are shown in Fig. 6.
@@ VI-B Implementation
Both C3 and the OSC were implemented in C++ with the help of the Drake robotics library [38]. Both controllers, as well as franka ros and the LCM to ROS bridges, are run on the same desktop with a Intel i7-8700K processor. The QP step of C3 was solved using OSQP [35], while the MIQP projection was solved using Gurobi [14]. The OSC QP was solved using OSQP [35] at 1000 Hz. We tune the OSC and C3 parameters by executing the task in the Drake [38] simulator using the hydroelastic contact model [28]. We directly apply the parameters that were tuned in simulation on hardware without additional tuning.
@@ VII Results
We performed multiple experiments to validate the robustness and generality of our framework. First we ablate our design decision to include the force tracking objective in the OSC by running experiments with and without that objective. We then demonstrate the reliability by continuously executing the experiment without manual resetting. Then, we demonstrate the robustness of our method to inaccurate models of mass and inertia by placing objects on the tray. We use the same controller parameters for all three targets and across all the demonstrations, and footage of the experiments can be seen in the supplemental video.
@@ VII-A Force Tracking Ablation
First we ablate the contribution of tracking the end effector force by executing 10 experiments with and without the tracking objective. The tracking controller with the end effector force objective succeeded for 80% of the trials, failing once when trying to reach the second target when lifting an unbalanced tray and failing once to reach the third target when the tray slipped off in the direction of the robot base. The trajectories of the end effector and tray for an execution are shown in Fig. 8 and Fig. 10. The tracking controller without the end effector force objective succeeded for 30% of the trials, failing seven times to reach the third target due to the tray either colliding when the supports or slipping off in the direction of the robot base.
@@ VII-B Reliability Test
The reliability of our method is evaluated by repeatedly executing the task without intervention. This is possible without manual resetting because the final target position coincides with the initial position, and thus we treat tracking error from the previous execution as unstructured perturbations to the initial state. Our method was able to complete six full cycles before failing due to the tray reach the position threshold (within 5cm) of the third target.
@@ VIII Limitations
While our controller is fairly robust to mass and inertia, it is not robust to the height of the supports. This is not surprising as the tray, supports, and robot are stiff, leading to senstivity at the boundary between contact and no-contact. However, this can be addressed by quickly adapting the LCS parameters [22] or generating the contact geometry and system dynamics entirely [20] [30].
Another limitation of this method originates from the LCS model used by the MPC. A known limitation of the LCS model is that it only is a linear representation of the contact decisions, meaning that it cannot consider contacts beyond the contact boundary. For example, if the end effector moves entirely out from underneath the tray, the gap function $\phi$ will shift and instead consider only horizontal contact between the end effector and tray. In this scenario, the MPC cannot find a solution for the task as it cannot reason about a path to go back under the tray. This linear representation of contact boundaries can be limiting when applying this method to certain dynamic manipulation tasks such as flipping or scooping, however this can be addressed by linearizing about a reference trajectory with more informative configurations. On the other hand, dynamic manipulation tasks such as batting and throwing should, in principle, fit well within our method.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1845298. Additional funding was provided from the AI Institute.